### 06 May 2026 v3

# Information Accessibility Limits in Structured NP Search

Abstract
We study the problem of locating violating principal minors in structured matrix families that lie near the boundary of P-matrices and admit sparse violations under perturbation. Viewing violation search as an information acquisition problem, we show that, despite strong underlying structure, the location of a violation may be globally encoded and not accessible through local queries under a restricted interaction model.
This leads to an information-theoretic bottleneck: each query reveals only vanishing information about the violating subset, so that polynomially many queries accumulate insufficient information to identify it. Using mutual information and Fano’s inequality, we show that any algorithm restricted to polynomially many queries cannot recover the violating subset with constant success probability.
Our analysis highlights a distinction between structure and accessibility: even highly structured problems can be computationally intractable when the information required to locate a solution is not accessible through the available queries.
Keywords: NP search; P-matrix; information theory; mutual information; information accessibility; information-theoretic lower bounds; Fano’s inequality.
1 Introduction
The study of P-matrices - matrices whose principal minors are all positive - plays a central role in optimization, complementarity problems, and numerical analysis. A fundamental computational question is to determine whether a given matrix is a P-matrix. Coxson [1] showed that this decision problem is co-NP-complete, suggesting that the associated search problem of locating a violating principal minor may also be computationally difficult.
In this work, we study this search problem. While the decision problem asks whether a violation exists, the search problem requires identifying its location among an exponentially large family of principal submatrices. Rather than analyzing this difficulty purely through computational complexity, we adopt an information-theoretic perspective.
Our central thesis is that the difficulty of this search arises not from a lack of structure, but from limitations on information accessibility. Even when a matrix exhibits strong algebraic regularity, the information needed to isolate a violating minor may be globally encoded and inaccessible through local queries. In particular, search becomes infeasible under the given interaction model when the rate of information acquisition vanishes, regardless of available computational power.
To investigate this phenomenon, we introduce a structured family of rank-one perturbations of P-matrices:
$A(u,v)=M+uv^{\top},$ (1.1)
where $M\in\mathbb{R}^{n\times n}$ is a P-matrix and $u,v\in\mathbb{R}^{n}$. In an appropriate parameter regime, all principal minors remain positive, while small perturbations can introduce sparse violations without destroying the global structure.
We analyze the problem under restricted interaction models in which algorithms access the matrix only through local queries. This abstraction does not impose an artificial limitation, but rather makes explicit the constraints on information acquisition inherent in the problem. Under a suitable distribution over instances, polynomially many queries accumulate only vanishing mutual information about the location of the violation. Consequently, by Fano’s inequality, recovery with constant success probability is impossible using polynomially many queries within this model.
More broadly, our analysis yields an information-accessibility framework for search problems, in which complexity is governed by the relationship between the information required to specify a solution and the amount revealed by each interaction. When the latter is insufficient, reliable recovery cannot be efficiently achieved.
This perspective aligns with recent information-theoretic approaches to search, such as the psocid framework [3], which establishes limits under restricted access in structureless settings. The present work extends this viewpoint to structured problems, showing that limitations can arise when structural constraints do not translate into accessible information under the interaction model. In particular, the barrier is not the absence of structure, but the inaccessibility of the information that the structure encodes.
To relate these insights to randomized computation, we use Yao’s minimax principle [4] to connect average-case analysis with randomized algorithms under the same interaction model. Under this view, information-theoretic limitations derived for a distribution over instances inform the behavior of randomized algorithms within this framework.
Our results highlight a conceptual separation between structure and accessibility: a problem may possess rich internal structure, yet remain difficult if the information required to locate a solution is not accessible through the given interaction model.
This perspective is motivated more broadly by search problems arising in combinatorial settings, but we focus here on a structured matrix family to isolate the role of information accessibility. To this end, we adopt a uniform prior over the hidden witness, removing instance-specific structural bias and providing a clean baseline in which all information about the solution must be acquired through interaction.
Section 2 introduces the preliminaries. Section 3 presents the information-theoretic analysis. Section 4 interprets the results through indistinguishability. Section 5 concludes the paper.
2 Preliminaries
Let $M\in\mathbb{R}^{n\times n}$. For any subset $\alpha\subseteq[n]$, let $M_{\alpha}$ denote the principal submatrix indexed by $\alpha$. A matrix $M$ is called a P-matrix if
$\det(M_{\alpha})>0\quad\text{for all }\alpha\subseteq[n],\ \alpha\neq\emptyset.$
We consider the following search problem.
Violation Search Problem.
For a P-matrix $M$, let $A(u,v)$ be defined as in (1.1) with $u,v\in\mathbb{R}^{n}$. The problem is to find a subset $\alpha\subseteq[n]$ such that
$\det(A(u,v)_{\alpha})\leq 0,$
if such a subset exists.
This is the search counterpart of the co-NP-complete problem of recognizing P-matrices [1].
We are interested in regimes in which violations of the P-matrix property are sparse relative to the total number of principal minors. Let
$\mathcal{V}(A(u,v)):=\{\alpha\subseteq[n]:\det(A(u,v)_{\alpha})\leq 0\}$
denote the set of violating principal minors.
We consider instances where
$|\mathcal{V}(A(u,v))|\ll 2^{n},$
that is, only a small fraction of subsets violate the P-matrix condition.
A particularly informative regime is the single-violation regime, in which
$|\mathcal{V}(A(u,v))|=1.$
In this case, there exists a unique subset $w^{\star}$ such that
$\det(A(u,v)_{w^{\star}})\leq 0,$
while all other principal minors remain positive.
In this setting, the violation search problem reduces to identifying a single hidden subset $w^{\star}$ among $2^{n}-1$ candidates.
Remark.
Appendix A provides an example with $|\mathcal{V}(A(u,v))|=1$ for $n=6$, and Appendix B gives a construction achieving the single-violation regime for general P-matrices $M$.
In this regime, the violating subset lies near the P-matrix boundary: the global structure remains largely intact, while the violation is localized and weakly expressed in observable quantities. As a result, its effect on query outcomes is minimal, leading to limited information accessibility.
3 Information Accessibility and Search Complexity
We formalize our analysis through an information-theoretic perspective on search problems, which we refer to as the Information-Accessibility Framework.
Information-Accessibility Framework.
A search problem is specified by a hidden variable $W$ representing the solution, together with an interaction model that determines how an algorithm acquires information about $W$. The difficulty of the search is governed by two quantities:
  • the information requirement, measured by the entropy $H(W)$;
  • the information acquisition rate, determined by the information revealed through the interaction model.
The central question is whether the accessible information suffices to identify $W$.
3.1 Instantiation: P-Matrix Violation Search
We instantiate this framework in the setting of P-matrix violation search. For notational convenience, we write $N$ in place of $n$ in the remainder of the paper.
Let $A=A(u,v)$ be drawn from the perturbation model defined in (1.1). Define the set of violating subsets by
$\mathcal{V}(A):=\{\alpha\subseteq[N]:\det(A_{\alpha})\leq 0\}.$
In the single-violation regime, $|\mathcal{V}(A)|=1$, and we denote the unique violating subset by $W\subseteq[N]$.
We assume a distribution over instances such that the unique violating subset $W\in\mathcal{V}(A)$ is uniformly distributed over the admissible nonempty subsets of $[N]$. Hence,
$H(W)=\log_{2}(2^{N}-1)=\Theta(N),$
where entropy is measured in bits.
Although $A(u,v)$ is algebraically structured, the condition defining $\mathcal{V}(A)$ depends on coordinated global interactions among its entries, and the location of $W$ is not directly revealed by local properties of $A(u,v)$.
Induced oracle model.
We consider an induced oracle model in which the algorithm interacts with the input instance through a sequence of queries. This interaction can be viewed as a communication channel through which information about the solution is revealed.
To formalize information accessibility, the algorithm issues queries of the form: for a subset $\alpha\subseteq[N]$, observe
$Y:=[\,\det(A(u,v)_{\alpha})\leq 0\,]\in\{0,1\}.$
The algorithm may adaptively select queries based on past observations. Formally, at step $t$, the query $\alpha_{t}$ is a function of the transcript $T_{t-1}=(Y_{1},\ldots,Y_{t-1})$, and the response is
$Y_{t}=\bigl[\,\det(A(u,v)_{\alpha_{t}})\leq 0\,\bigr].$
In the single-violation regime, this oracle satisfies
$Y=1\quad\Longleftrightarrow\quad\alpha=W,$
so the interaction reduces to equality testing.
We emphasize that the algorithm does not have direct access to the entries of $A(u,v)$ or its full structure, but only to the binary responses returned by the oracle.
Let $w^{\star}$ denote the realization of $W$. We allow $p(N)$ parallel queries per round, where $p(N)$ is polynomially bounded, and let $T$ denote the (random) round of the first successful query. In the single-violation regime under the uniform prior, the determinant-sign query is equivalent to an equality probe for $W$. Hence, by the psocid framework [3], the expected first-hit time satisfies
$\mathbb{E}[T]=\Omega\!\left(\frac{2^{N}}{p(N)}\right).$
Remark (Uniform prior and Yao’s principle).
The uniform prior over $W$ represents maximal uncertainty about the location of the violating subset and ensures that $H(W)=\Theta(N)$. By Yao’s minimax principle [4], lower bounds established under this distribution extend to worst-case guarantees for randomized algorithms.
Remark (Scope of the model).
The induced oracle model abstracts the information accessible to the algorithm and does not assume direct inspection of all entries of $A(u,v)$. Accordingly, the lower bounds derived below apply to algorithms whose access to the instance is mediated through such queries.
3.2 Information-Accessibility Barrier
The interaction model limits the rate at which information about $W$ can be acquired. The following information-accessibility barrier, established in the psocid framework [3], applies to the oracle model considered here.
Theorem 3.1 (Information-accessibility barrier).
Under the induced oracle model and a uniform prior over the nonempty subsets $W\subseteq[N]$, for every polynomial $\mathrm{poly}(\cdot)$ and every constant $\delta>0$, no algorithm making at most $\mathrm{poly}(N)$ queries can recover $W$ with success probability at least $\delta$.
More precisely, for any (possibly adaptive and randomized) algorithm making at most polynomially many queries, let $\mathcal{F}$ denote the resulting random transcript. Then the total accumulated mutual information between $W$ and $\mathcal{F}$ satisfies
$I(W;\mathcal{F})=o(1).$
Proof sketch.
Consider any (possibly adaptive) sequence of queries. We allow $p(N)$ parallel queries per round, where $p(N)$ is polynomially bounded. Let $Y_{t}=(Y_{t,1},\dots,Y_{t,m})$ denote the vector of responses in round $t$, where $m\leq p(N)$.
All query outcomes are flattened into a single sequence of $q=\mathrm{poly}(N)$ scalar responses:
$\mathcal{F}_{q}:=(y_{1},\dots,y_{q}),$ (3.1)
where $\{y_{k}\}_{k\geq 1}$ enumerates the coordinates $Y_{t,j}$ in lexicographic order of $(t,j)$. This is purely a notational device: each outcome $Y_{t,j}$ appears exactly once in $\mathcal{F}_{q}$.
In the single-violation regime, each query $\alpha$ returns
$Y=[\alpha=W].$
Under the uniform prior over the nonempty subsets of $[N]$, and conditioned on any transcript of k1k-1 failed queries, the posterior distribution of $W$ is uniform over the remaining candidates. Hence, for the $k$-th query,
$\Pr(y_{k}=1\mid y_{<k})\leq\frac{1}{2^{N}-1-(k-1)}.$
Since $k\leq q=\mathrm{poly}(N)$, this probability is exponentially small. Thus each observation is Bernoulli with parameter
$p_{k}\leq\frac{1}{2^{N}-O(q)},$ (3.2)
and its binary entropy satisfies
$H(y_{k}\mid y_{<k})=h_{2}(p_{k})=O\!\left(p_{k}\log\frac{1}{p_{k}}\right)=O\!\left(\frac{N}{2^{N}}\right).$
By the chain rule,
$I(W;\mathcal{F}_{q})=\sum_{k=1}^{q}I(W;y_{k}\mid y_{<k})\leq\sum_{k=1}^{q}H(y_{k}\mid y_{<k})=O\!\left(\frac{qN}{2^{N}}\right)=o(1).$
Since $H(W)=\Theta(N)$, Fano’s inequality implies that no algorithm making polynomially many queries can recover $W$ with constant success probability. ∎
Corollary 3.2 (Impossibility of polynomial-time recovery in the induced oracle model).
Any polynomial-time algorithm making at most polynomially many queries cannot recover $W$ with success probability bounded away from zero.
Corollary 3.3 (Worst-case hardness via Yao [4]).
For every randomized algorithm using polynomially many queries, there exists an input instance on which recovery fails with constant probability.
Interpretation of the barrier.
Theorem 3.1 identifies information accessibility - not computational effort - as the fundamental bottleneck. Even with unlimited computation, the interaction model reveals information about $W$ at a vanishing rate, so polynomially many queries cannot bridge the gap between the $\Theta(N)$-bit uncertainty in $W$ and the $o(1)$ information obtainable under polynomial interaction; by Fano’s inequality, recovery with constant success probability is impossible using polynomially many queries under this interaction model.
Implication.
The above barrier applies to the induced oracle model for the structured matrix family considered here. It shows that the algebraic structure of $A(u,v))$ does not translate into accessible information under this interaction model. The difficulty therefore arises not from lack of structure, but from limitations on how information about the solution can be obtained.
3.3 Mechanism: Information Masking and Intrinsic Limitation
We now describe the mechanism underlying the information-accessibility barrier and relate it to the rank-one construction of sparse-violation instances.
Information masking.
Each principal minor $\det(A(u,v)_{\alpha})$ depends on coordinated global interactions among many entries of $A(u,v)$. A local perturbation affects many minors simultaneously and does not isolate the violating subset.
We refer to information masking as the phenomenon in which the signal defining the solution is distributed across many components of the instance, so that individual queries provide only coarse and non-local information. In particular, the signal associated with a violating subset is not localized but spread across an exponential family of candidate subsets.
Remark (Absence of local ordering structure).
The ordering of principal minors is not monotone with respect to set inclusion: a subset attaining the minimal determinant need not have subsets with smaller determinants, nor supersets with larger ones. Consequently, comparisons among nested or nearby subsets do not provide reliable information about which subset minimizes the determinant.
In general, the search problem considered here is to identify a subset $\alpha$ such that $\det(A(u,v)_{\alpha})\leq 0$, rather than to locate a subset attaining the minimal determinant. These two tasks coincide only in the single-violation regime, where the unique violating subset must also minimize the determinant. In this regime, however, the absence of local ordering implies that identifying the relevant subset requires aggregating information across a global family of candidates, rather than relying on local structural cues.
Intrinsic information limitation.
One might attribute the difficulty of detecting a negative principal minor in the near-boundary regime to numerical precision. However, our analysis assumes exact arithmetic and abstracts away finite-precision effects.
The rank-one construction clarifies that the difficulty lies not in creating a violation - a small perturbation suffices once an appropriate subset is known - but in acquiring the information needed to locate such a subset.
The information-accessibility barrier is therefore intrinsic: even under idealized computation, recovery is limited by the information revealed through queries rather than by computational power or numerical precision.
4 Statistical Indistinguishability and Information Accessibility
The information-accessibility barrier can be understood through the statistical behavior of the response transcript.
We work in the single-violation regime, where each instance has a unique violating subset $w$. For any (possibly adaptive) sequence of queries, let $\mathcal{F}_{q}:=(y_{1},\dots,y_{q})$ denote the response transcript defined in (3.1), where $q=\mathrm{poly}(N)$.
Under the uniform prior over the nonempty subsets of $[N]$, by (3.2) and the union bound,
$\Pr(\exists k\leq q:y_{k}=1)\leq\sum_{k=1}^{q}p_{k}\leq\frac{q}{2^{N}-O(q)}=o(1).$
Hence, with probability 1o(1)1-o(1),
$\mathcal{F}_{q}=(0,0,\dots,0).$
In this event, the transcript is identical for all realizations of $W$. Consequently, for any $w\neq w^{\prime}$, the corresponding response distributions $P_{w}$ and $P_{w^{\prime}}$ both assign probability 1o(1)1-o(1) to the same outcome. Thus, the family $\{P_{w}\}_{w}$ collapses to a common distribution, and the response transcript takes essentially the same value for all instances.
Therefore, the response transcript is asymptotically independent of $W$. Since $\mathcal{F}_{q}$ is almost deterministic, it carries only vanishing information about the identity of the violating subset. Formally,
$I(W;\mathcal{F}_{q})=o(1).$
Consequently, polynomially many queries cannot identify the violating subset: specifying $W$ requires $\Theta(N)$ bits, while the response transcript reveals only a vanishing amount of information. This recovers the same conclusion as the Fano-based argument.
Interpretation. Under the oracle interaction model, transcript distributions corresponding to different instances are statistically indistinguishable except on exponentially rare events. Consequently, the observable signal distinguishing different instances is extremely weak.
In the near-boundary regime, small perturbations of the parameters may switch the matrix between feasibility and violation while affecting observable quantities only slightly. As a result, transcript distributions induced by feasible and violating instances may remain nearly indistinguishable under restricted interaction, contributing to the resulting information-accessibility barrier.
This suggests that the weakness of the observable signal may not be solely an artifact of the oracle model, but may reflect a property of the instances themselves: the relevant information is globally encoded while only weakly expressed through observable quantities.
Conjecture. There exist distributions over instances of the form $A(u,v)$ under which finding some violating subset $\alpha$ satisfying
$\det(A(u,v)_{\alpha})\leq 0$
can be achieved with non-negligible probability, while reliably recovering a specific violating subset $w^{\star}$ (for example, one satisfying
$\det(A(u,v)_{w^{\star}})\approx-\epsilon$
for a prescribed $\epsilon>0$) remains computationally hard on average.
This gap is not captured by the oracle lower bound established in this work, and suggests a stronger form of information inaccessibility: the solution is globally encoded while the resulting observations remain statistically indistinguishable.
This conjecture suggests that the information-accessibility barrier may extend beyond oracle models, pointing toward a broader class of structured problems in which global encoding and limited observability jointly induce computational difficulty.
5 Concluding Remarks
We develop an information-theoretic perspective on locating violating principal minors in structured matrix families, where violations arise from coordinated global interactions and cannot be efficiently localized through individual queries.
In the sparse-violation regime, particularly when $|\mathcal{V}(A)|=1$, the location of the violation carries high entropy while distinguishing observations are exponentially rare. Near the boundary between feasibility and violation, query outcomes are dominated by identical negative responses, rendering instances statistically indistinguishable. Consequently, each query reveals only negligible information and provides no directional signal, making the search problem ill-conditioned from an information-theoretic perspective.
These observations suggest a refined principle: structured problems may exhibit information-theoretic barriers similar to structureless search when the relevant information, though present, is not accessible under the given interaction model. In such settings, computational difficulty is governed not by the absence of structure, but by limitations on information acquisition.
A key limitation of the present analysis is the use of a uniform prior and a simplified access model. In realistic settings, such as 3-SAT, clause overlap and adaptive inference introduce complex dependencies that may increase the effective rate of information acquisition. Algorithms such as PPSZ [2] demonstrate that these structural features can be exploited to improve search performance; however, their behavior also suggests that information extraction remains limited and must be accumulated over many steps.
From the information-accessibility perspective, reductions from globally encoded search problems, such as P-matrix violation, to clause-based formulations such as 3-SAT may be viewed as mechanisms that transform weakly accessible global structure into locally exploitable logical structure. The auxiliary variables introduced by such reductions may play a role not only in representation, but also in exposing information in a form accessible to inference procedures such as clause propagation, which can aggregate information across multiple local interactions.
Understanding how structural properties enable higher rates of information extraction, and whether fundamental limits exist within specific interaction models, remains an important direction for future research. More broadly, a central open problem is to characterize the regimes in which information-accessibility barriers arise and to quantify the limits on information flow imposed by problem structure.
A further direction is to study the information-acquisition dynamics of structure-exploiting algorithms such as PPSZ. One may conjecture that such algorithms exhibit distinct phases: an initial phase in which structural features (e.g., critical clauses) enable relatively effective information extraction, followed by a regime in which the remaining uncertainty becomes less accessible, leading to a reduced effective information-acquisition rate. Importantly, this transition need not reflect a loss of structural constraints, such as the existence of critical clauses, but rather a reduction in their accessibility under the algorithmic interaction.
It is natural to ask whether this behavior reflects an intrinsic property of structured search under the given interaction model.
Appendix Appendix A A $6\times 6$ Illustrative Example
In this appendix, we present a concrete example illustrating the perturbation model introduced in the main text. The example exhibits a single-violation regime, in which a structured matrix admits exactly one non-positive principal minor under a rank-one perturbation.
Base P-matrix.
Consider the matrix
$M=\begin{pmatrix}2.3&-4.8&-1.9&0&0&1.3\\ 2.6&3.4&3.8&0&0&2.6\\ -2.6&4.8&6.7&-1.1&1.3&-1.3\\ 1.3&2.4&-3.8&5.4&0&-1.3\\ -2.6&4.8&0&-1.1&4.9&-1.3\\ 0&-4.8&-1.9&0&-2.6&7.5\end{pmatrix}$.
Numerical verification shows that all $63$ nonempty principal minors of $M$ are positive. In particular, the smallest principal minor is approximately $0.272$, attained at the subset $\{2,3,4\}$. Thus, $M$ is a P-matrix.
Perturbation.
Let
[…]
Violation.
For this choice of $u$ and $v$, numerical evaluation shows that exactly one principal minor of $A(u,v)$, defined in (1.1), is non-positive. Specifically, the principal submatrix indexed by $w^{\star}=\{1,5\}$ satisfies
$\det\bigl(A(u,v)_{w^{\star}}\bigr)\approx-0.001.$
All other nonempty principal minors of $A(u,v)$ remain positive. For comparison, $\det\bigl(M_{w^{\star}}\bigr)=11.27,$ which is substantially larger than the minimal principal minor, $0.272$.
Remark.
This example demonstrates that a highly structured matrix, obtained via a smooth rank-one perturbation of a P-matrix, can exhibit a sparse violation of the P-matrix property.
In this instance, the violating subset is not clearly reflected in the values of nearby principal minors: most minors, including those of comparable or larger size, remain positive and stable. Thus, local evaluations of principal minors provide only limited information about the location of the violation.
This behavior contrasts with typical constraint-based search problems, such as 3-SAT, where local constraints often provide informative signals that help guide the search toward a solution. In the present setting, however, no obvious local pattern among principal minors indicates the location of the violation.
This observation is consistent with the information-masking phenomenon discussed in the main text: although the violation is uniquely determined, the associated signal appears to be distributed globally and is not readily recoverable from local structure.
The construction is obtained numerically and serves to illustrate the existence of regimes in which violations are sparse and information about their location is not readily accessible from local observations.
Appendix Appendix B Rank-One Construction of Sparse Violations in P-Matrices
We describe a procedure for constructing instances in the sparse-violation regime, where $|\mathcal{V}(A(u,v))|=1$. The key idea is to decrease the smallest principal minor via a rank-one perturbation, without targeting any fixed subset.
For a P-matrix $M\in\mathbb{R}^{n\times n}$, define
$f_{M}:=\min_{\alpha\subseteq[n],\,\alpha\neq\emptyset}\det(M_{\alpha}),\qquad\alpha^{\star}\in\arg\min_{\alpha}\det(M_{\alpha}).$
Thus $f_{M}$ is the smallest principal minor of $M$, and $\alpha^{\star}$ is a corresponding minimizing subset.
Algorithm 1 (Rank-one construction).
  1. 1.
    Compute $f_{M}$ and $\alpha^{\star}$. Choose a vector $u\in\mathbb{R}^{n}$ such that
    $u_{i}>0\text{ for }i\in\alpha^{\star},\qquad u_{i}=0\text{ otherwise}.$
  2. 2.
    Define $\hat{v}\in\mathbb{R}^{n}$ supported on $\alpha^{\star}$, and choose its signs so that
    $\hat{v}_{\alpha^{\star}}^{\top}M_{\alpha^{\star}}^{-1}u_{\alpha^{\star}}<0.$
  3. 3.
    Consider the one-parameter family
    $A(\lambda):=M+\lambda u\hat{v}^{\top},\qquad\lambda>0.$
    By the matrix determinant lemma, for the subset $\alpha^{\star}$,
    $\det(A(\lambda)_{\alpha^{\star}})=\det(M_{\alpha^{\star}})\left(1+\lambda\,\hat{v}_{\alpha^{\star}}^{\top}M_{\alpha^{\star}}^{-1}u_{\alpha^{\star}}\right).$
  4. 4.
    Choose $\lambda>0$ such that
    $\det(A(\lambda)_{\alpha^{\star}})=-\epsilon,$
    where $\epsilon>0$ is small.
  5. 5.
    If necessary, reduce $\lambda$ slightly to ensure that there exists a unique subset $w^{\star}$ such that
    $\det(A(\lambda)_{w^{\star}})<0,$
    while all other principal minors remain positive. Note that $w^{\star}$ need not coincide with $\alpha^{\star}$.
  6. 6.
    Set $v:=\lambda\hat{v}$ and define $A(u,v):=M+uv^{\top}$.
Upon termination, the matrix $A(u,v)$ satisfies
$\det(A(u,v)_{w^{\star}})<0,$
while all other principal minors remain positive, yielding a single-violation instance.
Remarks.
  • In favorable cases, or under sufficiently small perturbations, the construction yields $|\mathcal{V}(A(u,v))|=1$.
  • Although the construction is initialized using a specific subset $\alpha^{\star}$, identifying the smallest principal minor requires solving a global minimization over all subsets. This involves the minimum over exponentially many subsets and is therefore generally non-differentiable. As a result, the identity of the minimizing subset may change discontinuously under perturbations.
  • The rank-one perturbation is applied only after identifying the globally minimal subset $\alpha^{\star}$. Thus, the difficulty of constructing a single-violation instance $A(u,v)$ lies in locating $\alpha^{\star}$ via a global search, rather than in the subsequent rank-one perturbation itself.
  • Moreover, the magnitude of the violating principal minor can be made arbitrarily small, making it increasingly difficult to distinguish it from nearby positive minors.
  • The example in Appendix A employs a different selection of $u$ and $v$ from Algorithm 1, illustrating that valid constructions can arise from a much broader choice of perturbation directions.